Company XYZ purchased equipment and gave a three-year note with maturity value of $12,006. The annual discount rate for the note was 14%, if XYZ has to make equal payments at the end of each quarter for three years (quarterly discounting), Then the payment each time should be $827.87. True or False
Answer #1

FV =12006
Number of Periods =3*4 =12
Rate per quarter =14%/4 =3.5%
The Payment per quarter =FV/(((1+r)^n-1)/r)=12006/(((1+3.5%)^4-1)/3.5%)=2848.44

​​​​​​​Hence the statement is false.